# spacing=nonfrench

### v2

@@ 6.2 The MeshGEMV algorithm
    Aggregation: The aggregation step primarily involves using the two-way K-tree allreduce we design. The key steps as follows: (i) In the 1st-phase, each group performs group reduction and obtains the partial sum of $C_{sub}$ at the root core of each group. (ii) In the $k$th-phase, the results from the $(k-1)$ th-phase are reduced to the root cores of each group in the $k$th-phase. After K times repeating, $C$ can be obtained by concatenating the $C_{sub}$ from all K-tree root cores. (iii)  Optionally, a broadcast operation from the root core of the $K$-tree may follow, depending on whether continuous GEMV is required.
Scalability Analysis. As shown in 1 of Figure 8, this method scales efficiently with parallelism and meets the L property by selecting an appropriate $K$. It requires $K+1$ paths at the tree root core but allows flexible adjustment of $K$ to address R based on hardware limitations.
However, a larger $K$ is not always better, as it depends on $N$ and R constraints. Additionally, larger $K$ increases routing complexity and overhead. Considering these factors, we have chosen $K=2$ for our current implementation evaluated in the following sections.
@@ 7 Evaluation
    WaferLLM’s MeshGEMM and MeshGEMV achieve strong performance and scalability over state-of-the-arts (§7.2);
@@
    WaferLLM’s shift-based KV cache management enables over 360$\times$ more token capacity (§7.4);
@@
    WaferLLM on Cerebras WSE-2 achieves up to 38.6$\times$ throughput and 1.7$\times$ energy efficiency compared to vLLM on A100 in LLM inference (§7.5).
Experiment setup. We evaluate WaferLLM on a server with Cerebras WSE-2. WSE-2 has 850,000 Cores, each with a Compute Engine (CE) operating at a maximum 1.1 GHz. Each clock cycle can fetch two 32-bit operands from SRAM, perform a multiply-accumulate operation, and then write back to SRAM. Each core also has a fabric router that can send or receive 32-bit messages from neighbouring cores with a single clock cycle. Additionally, each core contains 48KB of SRAM, with the chip totalling 40GB of aggregated SRAM [23].
@@
LLM models. Our evaluation includes various representative LLMs of different sizes and architectures. Specifically, LLaMA3-8B and LLaMA2-13B are widely used open-source LLMs, with LLaMA3 using group-query attention instead of multi-head attention to reduce KV cache usage. CodeLLaMA-34B is a specialized LLM for coding tasks, while QWen2-72B, another popular LLM, is renowned for its high model quality.
@@ 7.1 LLM inference
We first report the end-to-end performance of WaferLLM compared to T10 and Ladder. To provide deeper insights, we further analyze the performance by breaking down the execution into prefill and decode phases.
Model Seqlen In/Out 2048/128 4096/128 2048/2048 4096/4096
LLaMA3-8B WaferLLM 764.4 604.38 2370.33 2480.4
@@
Ladder 1.18 1.05 7.4 8.72
LLaMA2-13B WaferLLM 473.9 413.98 1690.28 1848
T10 2.6 2.51 35 58.27
Ladder 0.7 0.69 4.93 6.14
Table 2: End-to-end LLM inference throughput (tokens/s)
End-to-end throughput. Table 2 shows the inference throughput (i.e., tokens per second) of LLaMA3-8B and LLaMA2-13B on different configurations of input and output sequence length. WaferLLM uses core configurations optimized for the best performance with each model. In LLaMA3-8B, we use 660$\times$660 cores for prefill and 360$\times$360 for decode. In LLaMA2-13B, we use 750$\times$750 cores for prefill and 375$\times$375 for decode. CodeLLaMA-34B and QWen-72B are not included due to the memory constraint of WSE-2.
Compared to T10, WaferLLM achieves 160$\times$ speedup on average, up to 180$\times$, for short sequence generation tasks such as 4096 and 2048 input context lengths with 128 tokens output. For longer tasks, with input context lengths of 4096 and 2048 tokens and output lengths of 4096 and 2048 tokens, WaferLLM achieves 36$\times$ on average and up to 48$\times$. Although T10 designs the compute-shift model that considers the memory constraints (M) and routing resource limits (R) of a PLMR device, it does not account for the cores interconnected by a mesh NoC. thus failing to address varying hop distances (L) and scale to millions of cores (P), highlighting the need for new system designs in massive-scale NUMA architectures.
@@
Model Core Config 480$\times$480 600$\times$600 720$\times$720
LLaMA3-8B WaferLLM 20320.6 25037.22 27686.45
T10 175.01 156.62 132.82
Ladder 61.82 42.31 31.32
LLaMA2-13B WaferLLM 13685.10 16854.21 17498.28
T10 121.02 100.53 81.28
Ladder 47.25 33.14 24.23
CodeLLaMA-34B WaferLLM 5471.43 7540.13 8526
T10 49.06 46.77 41.23
Ladder 30.01 23.14 17.67
QWen2-72B WaferLLM 2785.19 3775.53 4421.58
T10 24.89 23.48 21.50
Ladder 16.77 12.80 10.12
Table 3: Prefill throughput (tokens/s)
@@
Prefill throughput. Table 3 shows the prefill throughput (i.e., input tokens processed per second) for an input sequence length of 4096, using core configurations from 480$\times$480 to 720$\times$720. For CodeLLaMA-34B and QWen2-72B, which exceed the memory capacity of WSE-2, we evaluate a subset of layers and scale the results proportionally due to their uniform layer structure.
@@
WaferLLM achieves significant speedups over T10 and Ladder by effectively addressing all PLMR properties. As discussed in §2, GEMM is the primary bottleneck, and MeshGEMM substantially enhances WaferLLM ’s prefill performance, analyzed in detail in §7.2.
@@
Additionally, WaferLLM scales throughput with increasing cores across all models. For instance, WaferLLM achieves a 1.6$\times$ scaleup on QWen2-72B and a 1.4$\times$ scaleup on LLaMA3-8B when scaling from 480$\times$480 to 720$\times$720 cores. This improved scalability is primarily because larger models better utilize the device’s compute resources, reducing the relative impact of NoC communication. In contrast, T10 and Ladder fail to scale effectively, with throughput even declining as more cores are added, due to persistent communication latency bottlenecks.
Model Core Config 420$\times$420 540$\times$540 660$\times$660
LLaMA3-8B WaferLLM 2699.94 2501.54 2243.25
T10 418.27 339.43 265.12
Ladder 14.6 13.09 11.42
LLaMA2-13B WaferLLM 2039.22 1899.4 1739.78
T10 341.83 270.79 233.72
Ladder 11.01 9.93 9.07
CodeLLaMA-34B WaferLLM 1450.77 1407.68 1359.18
T10 278.24 222.41 222.41
Ladder 6.07 6.15 5.77
QWen2-72B WaferLLM 839.71 824.3 787.08
T10 168.5 132.97 114.56
Ladder 3.23 3.29 3.38
Table 4: Decode throughput (tokens/s)
@@
Decode throughput. Table 4 shows decode throughput for core configurations from 420$\times$420 to 660$\times$660. For CodeLLaMA-34B and QWen2-72B, we evaluate a subset of layers and scale the results.
@@
By addressing all PLMR properties, WaferLLM achieves an average speedup of 5.7$\times$ (up to 6.5$\times$) over T10 and 217$\times$ (up to 260$\times$) over Ladder with 420$\times$420 cores.
Unlike prefill, decode throughput may decrease with more cores due to increased NoC communication latency, which impacts GEMV performance. Additionally, decode’s 6.5X improvement over T10 is less than prefill’s 160X because it involves less data transfer, limiting WaferLLM ’s communication advantage.
@@ 7.2 MeshGEMM
Figure 9: MeshGEMM vs. SUMMA vs. Cannon
We compare MeshGEMM with Cannon [7] and SUMMA [39] across different core scales and matrix sizes.
Scaling the number of cores. Figure 9 shows MeshGEMM with a fixed matrix size while scaling the number of cores. MeshGEMM achieves the lowest execution cycles by leveraging the INTERLEAVE operation to minimize communication overhead. It demonstrates stronger scalability, maintaining over 70% computational efficiency even near the hardware limit. In contrast, SUMMA and Cannon exhibit poor scalability, with computational efficiency falling below 50% with 720$\times$720 cores, primarily due to high communication overhead.
Additionally, increasing the cores does not consistently yield benefits. In GEMM 2K, scaling cores from 540$\times$540 to 720$\times$720 demonstrates diminishing returns: while computational cycles decrease, the additional communication overhead from more rounds of shifting negates these benefits. As a result, the total execution cycles for Cannon and SUMMA worsen significantly. In contrast, MeshGEMM mitigates these effects effectively, allowing further core scaling if required to address other constraints, such as the need for larger aggregate memory to accommodate data.
An interesting observation in Figure 9 is that for GEMM 8K, communication cycles decrease as core count increases. This occurs because GEMM 8K processes large data volumes and has lengthy computations, allowing full overlap with communication. In this scenario, communication becomes bandwidth-bound rather than latency-bound, and increasing core count boosts aggregated network bandwidth, resolving the bottleneck.
Scaling matrix size. We also evaluate MeshGEMM with larger matrix sizes, transforming GEMM into a more compute-intensive operation. At large scales, though the cost of communication becomes less significant, MeshGEMM maintains its scalability and outperforms SUMMA and Cannon by a wide margin, reducing total cycles by around 17%.
@@ 7.3 MeshGEMV
Figure 10 shows MeshGEMV with a fixed matrix size across an increasing number of cores. Compared to the baseline method, MeshGEMV significantly reduces communication cycles due to the efficient Two-way K-tree AllReduce, saving communication time and improving overall execution time by up to 4.6$\times$.
As the number of cores increases, the communication cost for MeshGEMV increases only slightly. In contrast, the baseline method’s linear reduce steps become increasingly costly as the number of cores increases, resulting in substantial performance degradation with more cores.
Scaling matrix size. We also evaluate MeshGEMV with larger matrix sizes. At larger scales such as $16K$, MeshGEMV maintains great scalability, with total execution cycles continuing to decrease as more cores are added. This contrasts with the baseline method, where a clear transition is observed: the savings in computational cycles diminish with increasing cores, and eventually result in negative gains.
@@ 7.4 Shift-based KV cache management
Table 5: Maximum tokens in generation
We also compare the shift-based KV cache management with the concat-based KV cache management implemented in PagedAttention. We evaluate KV cache capacity on LLaMA3-8B and LLaMA2-13B using the same settings as the end-to-end inference evaluation in §7.1. Table 5 shows that WaferLLM ’s shift-based KV cache management supports 360$\times$ and 385$\times$ more tokens than the concat-based method for LLaMA3-8B and LLaMA2-13B, respectively. This improvement results from balanced core utilization and the resolution of skewed data issues achieved by the shift-based approach.

### v3

@@ 6.2 The MeshGEMV algorithm
    Aggregation: The aggregation step primarily involves using the K-tree allreduce we design. The key steps are as follows: (i) In the 1st-phase, each group performs group reduction and obtains the partial sum of $C_{sub}$ at the root core of each group. (ii) In the $k$th-phase, the results from the $(k-1)$ th-phase are reduced to the root cores of each group in the $k$th-phase. After K times repeating, $C$ can be obtained by concatenating the $C_{sub}$ from all K-tree root cores. (iii) Optionally, a broadcast operation from the root core of the $K$-tree may follow, depending on whether continuous GEMV is required.
Scalability Analysis. As shown in 3 of Figure 8, this method scales efficiently with parallelism and meets the L property by selecting an appropriate $K$. It requires $K+1$ paths at the tree root core but allows flexible adjustment of $K$ to address R based on hardware limitations. Compared to pipeline and ring all-reduce, K-tree all-reduce essentially builds pass-through paths between distant cores by consuming routing resources, thereby reducing routing latency.
However, a larger $K$ is not always better, as it depends on $N$ and R constraints. Additionally, larger $K$ increases routing complexity and overhead. Considering these factors, we have chosen $K=2$ for our current implementation, evaluated in the following sections.
@@ 7 Evaluation
    WaferLLM’s MeshGEMM and MeshGEMV perform and scale strongly over the state-of-the-art (§7.2);
@@
    WaferLLM’s shift-based KV cache management enables over 360-385$\times$ more token capacity (§7.4);
@@
    WaferLLM on Cerebras WSE-2 achieves 10-20$\times$ e2e speedup compared to the optimal performance of SGLang on A100 GPUs-cluster, and about 30-40$\times$ speedup compared to a single A100 GPU. WaferLLM also achieves approximately 2.5$\times$ better energy efficiency (§7.17.5).
Experiment setup. We evaluate WaferLLM on a server with Cerebras WSE-2. WSE-2 has 850,000 Cores, each with a Compute Engine (CE) operating at a maximum of 1.1 GHz. Each clock cycle can fetch two 32-bit operands from SRAM, perform a multiply-accumulate operation, and then write back to SRAM. Each core also has a fabric router that can send or receive 32-bit messages from neighbouring cores with a single clock cycle. Additionally, each core contains 48KB of SRAM, with the chip totalling 40GB of aggregated SRAM [24].
@@
We use the Nvidia A100 for GPU comparison experiments, which shares the same 7nm process as the WSE-2 for fairness. The experiments utilize up to 16 GPUs across two nodes (2$\times$8 configuration). Within a single node, the eight A100 GPUs are interconnected via NVLink, while inter-node communication is handled through a high-performance InfiniBand (IB) network. We use SGLang[53], one of the highest-performing LLM inference systems for GPU-based experiments.
Experiment metric. To evaluate the maximum per-request throughput performance of LLM inference systems on different hardware, we define Throughput per Request (TPR) as a key metric. TPR is derived from the more widely used Time per Output Token (TPOT), with $\text{TPR}=\frac{1}{\text{TPOT}}$.
LLM models. Our evaluation includes various representative LLMs of different sizes and architectures. Specifically, LLaMA3-8B and LLaMA2-13B are widely used open-source LLMs, with LLaMA3 using group-query attention instead of multi-head attention to reduce KV cache usage. For LLaMA2-13B, we modified the model to remove the 4K context length limitation to evaluate system performance across different input and output lengths. Additionally, due to the tensor parallelism constraints, the number of attention heads of a model must be divisible by the number of GPUs used. We did not conduct the multi-GPU experiment of LLaMA-2 on a 16-GPU cluster. CodeLLaMA-34B is a specialized LLM for coding tasks, while QWen2-72B, another popular LLM, is renowned for its high model quality.
@@ 7.1 LLM inference
We first report the end-to-end performance of WaferLLM compared to T10 and Ladder on WSE-2 and SGLang on the A100 GPU cluster. Then we further analyze the performance to provide deeper insights by breaking down the execution into prefill and decode phases.
Model Device Input/Output Seqence Length
2048/128 4096/128 2048/2048 4096/4096
LLaMA3-8B WSE-2 WaferLLM 764.4 604.4 2370.3 2459.0
@@
Ladder 1.2 1.1 7.4 8.7
A100 GPU (SGLang) 1 34.8 31.1 36.5 78.4
8 117.2 109.0 128.4 256.1
2$\times$8 73.7 70.2 79.3 162.5
LLaMA2-13B WSE-2 WaferLLM 473.9 414 1690.3 1826.0
T10 2.6 2.5 35.0 58.3
Ladder 0.7 0.7 4.9 6.1
A100 GPU (SGLang) [^3] No 2$\times$8 GPUs, due to model architecture and tensor parallelism. 1 20.4 17.1 21.1 47.9
8 79.6 70.5 86.9 172.4
Table 2: End-to-end LLM inference TPR
End-to-end throughput. Table 2 shows the inference TPR of LLaMA3-8B and LLaMA2-13B on WSE-2 and A100 with different input and output sequence lengths. Here, the end-to-end throughput is calculated as the total number of tokens generated during the decode phase divided by the total time spent in the prefill and decode phases. WaferLLM uses core configurations optimized for the best performance with each model. In LLaMA3-8B, we use 660$\times$660 cores for prefill and 360$\times$360 for decode. In LLaMA2-13B, we use 750$\times$750 cores for prefill and 375$\times$375 for decode. CodeLLaMA-34B and QWen2-72B are not included due to the memory constraint of a single WSE-2 chip.
Compared to T10, WaferLLM achieves 160$\times$ speedup on average, up to 180$\times$, for short sequence generation tasks such as 4096 and 2048 input context lengths with 128 tokens output. For longer tasks, with input context lengths of 4096 and 2048 tokens and output lengths of 4096 and 2048 tokens, WaferLLM achieves 36$\times$ on average and up to 48$\times$. Although T10 designs the compute-shift model that considers the memory constraints (M) and routing resource limits (R) of a PLMR device, it does not account for the cores interconnected by a mesh NoC. Thus, failing to address varying hop distances (L) and scale to millions of cores (P), highlighting the need for new system designs in massive-scale NUMA architectures.
@@
Model Methods WSE-2 Cores # A100 GPUs (SGLang)
480$\times$480 600$\times$600 720$\times$720 1 8 2$\times$8
LLaMA3-8B WaferLLM 20320.6 25037.2 27686.5 13988.3 17361.6 13994.2
T10 175.0 156.6 132.8
Ladder 61.8 42.3 31.3
LLaMA2-13B WaferLLM 13685.1 16854.2 17498.3 7805.1 12287.1 /3
T10 121.3 100.6 81.3
Ladder 47.3 33.1 24.2
CodeLLaMA-34B WaferLLM 5471.4 7540.1 8526 5382.5 7155.5 6409.2
T10 49.1 46.8 41.2
Ladder 30.1 23.1 17.7
QWen2-72B WaferLLM 2785.2 3775.5 4421.6 1677.3 3803.8 3750.5
T10 24.9 23.5 21.5
Ladder 16.8 12.8 10.1
Table 3: Prefill Throughput per Request (TPR)
@@
Prefill throughput. Table 3 shows the prefill performance for an input sequence length of 4096, using core configurations from 480$\times$480 to 720$\times$720. For CodeLLaMA-34B and QWen2-72B, which exceed the memory capacity of WSE-2 and a single A100, we evaluate a subset of layers and scale the results proportionally due to their uniform layer structure.
@@
WaferLLM achieves significant speedups over T10 and Ladder by effectively addressing all PLMR properties, an average speedup of  160$\times$ (up to 178$\times$) over T10 and  270-450$\times$ over Ladder. As discussed in Section §2, GEMM is the primary bottleneck, and MeshGEMM substantially enhances WaferLLM’s prefill performance, analyzed in detail in Section §7.2.
@@
Additionally, WaferLLM scales throughput with increasing cores across all models. For instance, WaferLLM achieves a 1.6$\times$ scaleup on QWen2-72B and a 1.4$\times$ scaleup on LLaMA3-8B when scaling from 480$\times$480 to 720$\times$720 cores. In contrast, T10 and Ladder fail to scale effectively, with throughput even declining as more cores are added. This is mainly because T10 and Ladder do not account for the spatial locality characteristic (L) of highly uniformly distributed memory, which can lead to frequent long-distance data reads across cores and increased communication overhead.
Model Methods WSE-2 Cores # A100 GPUs (SGLang)
420$\times$420 540$\times$540 660$\times$660 1 8 2$\times$8
LLaMA3-8B WaferLLM 2699.9 2501.5 2243.3 78.9 260.4 164.6
T10 418.3 339.4 265.1
Ladder 14.6 13.1 11.4
LLaMA2-13B WaferLLM 2039.2 1899.4 1739.8 48.7 175.8 /3
T10 341.8 270.8 233.7
Ladder 11.0 9.9 9.0
CodeLLaMA-34B WaferLLM 1450.8 1407.7 1359.2 26.1 100.4 84.5
T10 278.2 222.4 193.1
Ladder 6.1 6.2 5.8
QWen2-72B WaferLLM 839.7 824.3 787.1 10.6 51.2 48.7
T10 168.5 133.0 114.6
Ladder 3.2 3.3 3.4
Table 4: Decode Throughput per Request (TPR)
@@
Decode throughput. Table 4 shows decode throughput for core configurations from 420$\times$420 to 660$\times$660. For CodeLLaMA-34B and QWen2-72B decode benchmark on WSE-2 and single A100, we evaluate a subset of layers and scale the results.
@@
By addressing all PLMR properties, WaferLLM achieves an average speedup of 5.7$\times$ (up to 6.5$\times$) over T10 and 217$\times$ (up to 260$\times$) over Ladder.
Compared to the 160$\times$ speedup WaferLLM achieves over T10 during prefill, it offers only about a 6.5$\times$ speedup during decode. However, WaferLLM maintains a consistent 200-500$\times$ advantage over Ladder in both prefill and decode stages. This is mainly because prefill involves dist-GEMM computations, which require each data element to traverse cores along the same row or column sequentially. In contrast, the dist-GEMV allreduce operation in decode only requires traversal across cores in the same row or column, without enforcing a strict access order. While T10’s compute-shift paradigm does not consider the spatial locality of cores on a mesh NoC, it still accounts for distributed memory, which allows performance gains when memory (communication) accesses are order-independent. In contrast, Ladder is entirely designed for a shared memory architecture, leading to severe communication overhead regardless of whether memory access order is required.
Comparison with SGLang. As shown in Table 2,3 and 4, WaferLLM demonstrates clear advantages over SGLang + A100 GPU clusters in prefill, decode tasks, and also the end-to-end performance across models ranging from 8B to 72B. More detailed comparisons with GPUs are discussed in Section §7.5.
It is worth noting that T10 and Ladder fail to leverage the powerful hardware capabilities of the WSE-2 when compared to SGLang. For example, in prefill tasks and for Ladder in decode tasks, their performance is even worse than that of SGLang running on a single A100. This highlights that designing high-performance systems and algorithms on wafer-scale chips must carefully consider the PLMR model; otherwise, not only will the massive compute potential of the chip go underutilized, but performance may even degrade due to the highly uniform distributed memory resulting from the large-scale mesh NoC.
@@ 7.2 MeshGEMM
Figure 9: MeshGEMM vs. SUMMA & Cannon
We compare MeshGEMM with Cannon [6] and SUMMA [42] on WSE-2 across different core scales and matrix sizes.
Scaling core count. Figure 9 shows that across matrices of different sizes, MeshGEMM can achieve the lowest latency than SUMMA and Cannon by scaling up the number of cores. Compared to the pipeline broadcast in SUMMA and head-to-tail transmission in Cannon, interleave transmission in MeshGEMM can minimize per-step communication overhead and, as much as possible, overlap communication with computation during large-scale fine-grained parallel execution. It demonstrates stronger scalability, maintaining over 70% computational efficiency even near the hardware limit. In contrast, SUMMA and Cannon exhibit poor scalability, with computational efficiency falling below 50% with 720$\times$720 cores, primarily due to the communication overhead in SUMMA and Cannon, increasing with the parallelism scale.
Additionally, increasing the number of cores does not always bring performance gains, especially for small matrix GEMM tasks. For example, in GEMM 2K, when scaling from 360$\times$360 to 720$\times$720 cores, the end-to-end latency of SUMMA and Cannon increases instead of decreasing. This is because the per-core computation cost drops sublinearly as cores increase beyond a threshold (due to fixed overheads like function calls and logic checks). At the same time, communication overhead grows linearly, as seen from the shaded areas in Figure 9, leading to worse overall performance. In contrast, MeshGEMM only shows a slight communication overhead increase at 720$\times$720 cores and maintains stable end-to-end latency. This is because its interleave communication bounds the per-step communication cost to a constant, independent of core count (the total communication overhead grows only because the number of steps increases in all dist-GEMM algorithms). Thus, MeshGEMM can achieve a better overlapping between communication and computation even under large-scale fine-grained parallelism (each core is only responsible for a small amount of data).
Scaling matrix size. We also evaluate MeshGEMM with larger matrix sizes, transforming GEMM into a more computationally intensive operation. At large scales, though the cost of communication becomes less significant, MeshGEMM maintains its scalability and outperforms SUMMA and Cannon by a wide margin, reducing total cycles by around 17%.
An interesting observation in Figure 9 is that for GEMM 8K, communication cycles decrease as core count increases. This occurs because when processing large data volumes, communication is bandwidth-bound rather than latency-bound. Increasing the number of cores not only boosts aggregate compute power to reduce computation overhead but also increases aggregate bandwidth to lower overall communication cost.
@@ 7.3 MeshGEMV
Communication latency bottleneck. Figure 10 shows that communication becomes the primary bottleneck for dist-GEMV computations on WSE-2, especially when the parallelism scale is large relative to the computation per core, where communication overhead can dominate 90% of the total. However, MeshGEMV significantly reduces communication overhead compared to the Cerebras baseline GEMV, achieving about 4.6$\times$ higher end-to-end performance. This is because large-scale dist-GEMV requires long-distance allreduce communication, and MeshGEMV uses K-tree Allreduce to maximize parallelism in the allreduce process, minimizing communication and computation costs along the critical path.
Scaling core count and matrix size. We also evaluate MeshGEMV with larger matrix sizes. For GEMV 8K and 16K compared to 2K, the baseline shows a trend that the end-to-end overhead first decreases and then increases. The initial decrease comes from increased core count, where the higher aggregate compute power and bandwidth reduce computation and communication costs. However, as the core count grows, the communication latency of allreduce eventually overtakes computation and bandwidth as the dominant bottleneck. In contrast, MeshGEMV’s inflection point appears later, which means that for the same data size, MeshGEMV’s communication overhead grows much more slowly with increasing core count compared to the baseline, demonstrating better scalability. This enables more flexible on-chip mapping policies, such as using more cores to store additional LLM parameters to meet memory limitations (M) without introducing significant extra overhead.
@@ 7.4 Shift-based KV cache management
Table 5: Maximum decode output length
We also compare the shift-based KV cache management with the concat-based KV cache management in PagedAttention. We evaluate KV cache capacity on LLaMA3-8B and LLaMA2-13B using the same settings as the end-to-end inference evaluation in Section §7.1. Table 5 shows that WaferLLM’s shift-based KV cache management supports 360$\times$ and 385$\times$ more tokens than the concat-based method for LLaMA3-8B and LLaMA2-13B, respectively. This improvement results from balanced core utilization and the resolution of skewed data issues achieved by the shift-based approach.